OpenStudy (anonymous):

simplify the expression 4^-3

OpenStudy (anonymous):

\[\Huge 4^{-3}=\dfrac{1}{4^3} \]

OpenStudy (unklerhaukus):

\[\frac{1}{4^3}=\frac{1}{4\times4\times4}\]